You have a 100-watt light bulb on your front porch that is turned on 10 hours every night. If the cost of electricity is 12 cents per KWHR (kilowatt hour), how much does that bulb add to your monthly electric bill?
1
Expert's answer
2019-04-24T09:57:59-0400

The electric bill amount=(work done)x(price)

"\rm{bill}=(\rm{power} \times \rm{time})\times(\rm{price})"

"=(0.1\:\rm{kW})\times (10\:\rm{hr}\times 30\:\rm{days})\times (12\:\rm{cents})=3.60\:\$"